# We Have a Package for You! A Comprehensive Analysis of Package Hallucinations by Code Generating LLMs

### v1

@@ 3 Research Questions
Adversary Model and Assumptions. In this work, we assume an adversary that wants to accomplish a package confusion attack by leveraging package hallucinations in the code generated by publicly-available commercial/closed-source (e.g., ChatGPT [5]) and open-source code generation LLMs (e.g., CodeLlama [71]). The target of the adversary are users who employ such LLMs for generating code snippets or entire programs, which they directly use in their codebase without any significant verification or modification. In other words, the target users fully trust the LLMs to include only valid package names in the generated code. We assume that the adversary has access to the same set of LLMs for code generation as the target users, and is unable to modify or manipulate the model and model parameters of these LLMs (e.g., via re-training or fine-tuning) before they are used by the victims. The adversary is able to determine a list of hallucinated packages in the code generated by these LLMs (for example, by cross-referencing on the package repository), and then to realize a package confusion attack the adversary is also able to create packages with the same names on the corresponding package repositories. These newly created packages corresponding to the fictitious packages generated by the LLM could contain malicious code or functionality to compromise the target users. Figure 1 provides a high-level overview of package hallucination exploitation process.
Research Questions. We now organize our investigation of package hallucinations by popular publicly-available LLMs into five broad Research Questions (RQ), as outlined below.
RQ1: How prevalent are package hallucinations while generating Python and JavaScript code using LLMs? By means of this RQ, our goal is to gain a comprehensive understanding of how often package hallucinations occur (i.e., rate of hallucinations) when using popular commercially available and open-source LLMs to generate Python and JavaScript code for a variety of different programming tasks.
RQ2: How are package hallucinations impacted by certain model settings? Here our goal is to comprehensively analyze how select model properties and settings, namely training data recency, model temperature, and decoding strategies impact the package hallucinations produced by the LLMs.
@@
RQ5: Is it possible to effectively mitigate package hallucinations using best practices in the literature and knowledge gained from earlier results? Through this RQ, we will investigate if code generating LLMs can be designed to reduce hallucinations. In this direction, we will comprehensively study if techniques such as Retrieval Augmented Generation (RAG), self-detected feedback, decoding strategies, and supervised fine-tuning are effective hallucination reduction strategies.
@@ 4 Experiment Design
To answer the RQs outlined above, we design several experiments to first iteratively prompt LLMs to generate code and then analyze the generated code. Our experimentation pipeline comprises of three distinct phases: (i) prompt dataset generation, (ii) code generation, and (iii) hallucination detection, each of which is described below:
@@ 4.1 Prompt Dataset
The experiments are designed to exhaustively test each LLM through a complete range of potential coding tasks. As existing datasets of coding prompts are used as benchmarks for correctness [9, 47] and contain only a limited number of prompts that lack diversity, we decided to create a new code prompt dataset for our experiments. Our goal was to create a dataset that was both realistic and representative of the coding task requests observed in practice by everyday users. To accomplish this, we employed two distinct approaches, as described below.
@@
Stack Overflow Dataset. In order to model the input prompts (to LLMs) in our experiments around real programmer questions, our first prompt dataset was created using Stack Overflow [75] questions across relevant programming topics and subject areas. Stack Overflow is a popular online question-and-answer service for software programmers and developers. To capture a wide range of topics, we utilized the “tag” feature of Stack Overflow, which allows users to label posts according to a subject matter. We included any tag that had over 5,000 questions and was also relevant to Python or JavaScript (the two programming languages we focus in this work, as detailed in Section 4.2). For each of the 240 tags that met this criteria, we extracted the 20 most up-voted questions, resulting in 4,800 prompts (i.e., 4,800 prompts for Python and 4,800 prompts for JavaScript).
@@
As more recent data is less likely to be included in the pre-training data of LLMs, we are also interested in investigating the correlation between data recency (i.e., how recently the question was asked on Stack Overflow) and model hallucination rate. To enable such an analysis, we ran two queries on Stack Overflow; one that only captured the most popular questions in the selected tags from 2023 and another that captured the most popular questions of all-time. By including the two different ranges of time, we effectively doubled the original number of prompts, for a total of 9,600 for each of the two languages. There is some overlap between the datasets for each language, as some of the tags pertain to general coding questions while others are language-specific. Also, not all questions asked on Stack Overflow may involve coding or require code to answer the question. For example, some questions might be about conceptual topics, such as software design principles or debugging strategies, which do not necessitate code snippets. Rather than attempting to filter out such prompts during the code generation phase, which is non-trivial and error-prone, the LLM is asked to answer the question and only provide code if necessary. In the end, this may result in a slightly lesser number of usable LLM-generated code samples, but is more realistic as LLMs are expected to accommodate imperfect user inputs.
@@
LLM-generated Dataset. As a majority of the programming tasks require some library/package, our next idea was to use the package repositories themselves as a good representation of the full spectrum of coding topics. For creating our second prompt dataset, we take the 5,000 most popular (according to number of downloads) Python and JavaScript packages and scrape the official package description as listed on pypi.org and npm.org, respectively. These descriptions are then input individually to the Llama-2 70B model with the following instruction: “Generate a prompt to produce Python/JavaScript code that would accomplish the same functionality as the following package description: [package description]”. This process worked remarkably well and generated roughly 4,800 prompts for Python and JavaScript each, resulting in two datasets of approximately the same size. Similar to the Stack Overflow dataset, we doubled the LLM-generated dataset for recency analysis by dividing it into two segments: packages most downloaded in the past year and packages most downloaded in the period before the past year (ensuring no overlap). When a package ranks among the most downloaded for both the past year and all-time, we give preference to the latter. During this process some packages with no description or descriptions in a non-English language were discarded. Both the Stack Overflow and LLM-generated datasets will be made publicly available through the project’s GitHub page. A truncated list from the LLM-generated dataset can be found in Appendix A.
@@ 4.2 Code Generation
Model Selection. For our experiments, we chose the top-ranked models from the EvalPlus leaderboard (as of January 2024) [47]. While creating our list, we disregarded the fine-tuned versions that were ranked below their corresponding foundational models, and only selected one fine-tuned version of the same foundational model of the same parameter size [47]. EvalPlus maintains a ranking of the top performing LLMs for code correctness according to a rigorous code synthesis evaluation framework. Our goal was to include a mix of top performing base models and a few of the best performing fine-tuned variants. We also included the GPT series of models (GPT-3.5, GPT-4, and GPT-4 Turbo) in our experiments, which currently hold the top rankings on the leaderboard. GPT models are widely considered as SOTA in terms of code generation models at the time of writing and add value to our experiments as representative commercial models. Table I provides a complete list of the models that we tested in our experiments.
TABLE I: Details of the models that were evaluated.
@@
Language Selection. In our experiments, we focus on two of the most popular programming languages, JavaScript and Python, according to GitHub’s 2023 Octoverse report [17]. The open-source package repositories for these languages, npm and PyPI, represent ecosystems of 4.5 million and 535 thousand packages, respectively [78]. Given the popularity of these two programming languages and the scale of these software ecosystems, they are more likely to be targets of adversaries and impacted by package hallucinations and related package confusion attacks. It is safe to assume that many open-source code generation models have been fine-tuned to program in these two specific languages, particularly Python due to its high popularity. Online services such as HuggingFace actively track the best coding models, specifically using Python and JavaScript as the two benchmark languages [47]. In total, we tested 16 different models (see Table I), fourteen of which were tested for both Python and JavaScript, while two fine-tuned Python-specific models, WizardCoder-Python and CodeLlama-Python, were only tested for Python.
Testing Environment. To create a uniform and fair testing environment, we utilize the text-generation-webui tool for LLMs, which supports a majority of open-source LLMs [62]. This tool provides both a GUI and an API that allows a user to use LLMs locally and offline. It also provides fine-grained control of parameters during testing which ensures that each model is evaluated using identical environments. We opted to use quantized versions of the open-source LLM models, using the GPTQ quantization method. Quantization is a process that reduces the precision of model parameters, improving inference speed and reducing memory requirements without significantly impacting performance. The GPTQ quantization method, specifically, uses a one-shot weight quantization method based on approximate second-order information that has a negligible effect on the accuracy of models, making it an ideal choice [14]. Additionally, quantized models better simulate the performance that a typical user can expect when running models on commercial grade hardware, making them more accessible and practical for everyday use. For testing uniformity, we use the same parameters and quantization precision for all open-source models, which are summarized in Appendix C. Model testing was conducted in two distinct computing environments - a Debian environment with 40 nodes, each equipped with 40 CPU cores, 1TB of RAM, and NVIDIA A100 or V100 GPUs and a Ubuntu system with 80 CPU cores, 750 GB of RAM, and 3 NVIDIA RTX 6000 GPUs. To generate code for our analysis, we query each of the LLMs (Table I) with prompts from the two datasets using the Python requests packages to interact with the text-generation-webui API. The message to the API contains a system message with specific instructions for the model and the prompt itself. An overview of the process, including system messages used during each step, are detailed in Appendix B. This results in a total of 19,200 code samples per model (16 Python tests + 14 JavaScript tests * 19,200 = 576,000 total code samples), which are further analyzed to determine which packages are required to execute the generated code.
@@ 4.3 Detection Methodology and Heuristics
Detection of hallucinated packages from LLM outputs or code samples is non-trivial. Before we outline our heuristics to detect packages in the generated Python and JavaScript code samples, let’s first discuss the notion of packages and modules in modern interpreted languages. To enable code modularity and reusability, interpreted programming languages such as Python and JavaScript allow for entities called modules. A module is a chunk of code, often in an external file, that performs a specific task or function. By encapsulating related code into modules, developers can organize their programs more efficiently and make them easier to maintain and share. A package is a collection of related modules that work together to provide certain functionality. To use a particular module in their source code, a developer must install the appropriate package into its development environment by first downloading the package from an online package manager or repository such as PyPI or npm, if it is not locally available, and then import the desired module into the code using appropriate import functions, e.g., import in Python or require in JavaScript.
These module names do not necessarily need to match the package names and the namespace for modules is not protected, i.e., different packages may include modules of the same name. This discrepancy poses a significant challenge for detecting package dependencies from the raw Python and JavaScript code, as the import/require statements that are typically included in code samples for importing modules do not have a unique mapping to package names. Thus, it would not be accurate to simply parse the generated code for import or require statements, as those statements refer to module names and not package names. To solve this problem, we employ the following three heuristics to determine/identify package names:
@@
Heuristic 2. For the second heuristic, each generated code sample is used as an input into the same model that generated it. The model is then asked (prompted) for a list of packages that would be required to run the given code. Our intuition is to mimic an actual user/developer who is using LLMs for code generation. If the user gets an error due to an uninstalled package when attempting to execute the generated code, they could query the model for the correct package to install. We wanted to replicate this intuitive process to identify package names required by the generated code.
Heuristic 3. As the third heuristic, we re-use the original prompt used to generate the code sample as an input to the model and ask the model to output package names that would be required to accomplish this coding task. Similar to the previous heuristic, this process of extracting package names simulates another approach users would take to obtain package names from the model that generated the code, if the required packages were not mentioned in explicit “pip install” and “npm install” commands. Once each model provides specific package names (through the three heuristics outlined above), we simply compare each package name to a master list of package names acquired from PyPI and npm, respectively (each list is as of 10 January, 2024). If a package name is not on the master list it is considered a hallucination. To measure LLMs’ propensity to produce hallucinated packages during code generation we use the package hallucination rate metric, which can be expressed as a simple ratio of the number of hallucinated packages to the total number of recommended packages.
@@ 5 Evaluation Results
In this section, we present the results of our experimental analysis related to RQ1 - RQ4. For RQs 2 through 4, we focus our analysis only on the Python programming language and a subset of the original models tested. Given the consistent results that we were able to obtain across both languages, this narrowed scope of discussion does not compromise the generalizability of the conclusions and allows for a deeper analysis of package hallucinations in a controlled setting. We selected GPT-4 Turbo, GPT-3.5, CodeLlama, and DeepSeek for in-depth analysis, representing the best-performing and most popular open-source models. The full set of hallucination results for all models, covering both Python and JavaScript, are presented in Appendix G.
@@ 5.1 Prevalence of Package Hallucinations (RQ1)
In our first experiment, our goal was to quantify the prevalence of package hallucinations across different models by generating and analyzing a large number of code samples. We conducted 30 tests (using 16 models for Python and 14 models for JavaScript, as outlined in Table I) producing a combined 576,000 code samples using both the Stack Overflow and LLM-generated datasets (Section 4.1). Each code sample was evaluated for hallucinations according to the heuristics defined in Section 4.3, which include parsing the generated code and prompting the model for packages twice per code sample, for a total of 1,152,000 package prompts across all tests.
@@
These 30 tests generated a total of 2.23 million packages in response to our prompts, of which 440,445 (19.7%) were determined to be hallucinations, including 205,474 unique non-existent packages (i.e. packages that do not exist in PyPI or npm repositories and were distinct entries in the hallucination count, irrespective of their multiple occurrences). GPT-4 Turbo resulted in the lowest overall hallucination rate at 3.59%, while DeepSeek 1B had the best hallucination rate among the open-source models at 13.63%. Python code resulted in fewer hallucinations than JavaScript (15.8% on average compared to 21.3% for JavaScript), a result that is expected, given that the JavaScript ecosystem contains nearly 10x the amount of packages compared to Python (npm contains 4.75 million while PyPI contains 548,000). Despite the difference in the hallucination rate between the two languages, there is a linear relationship between the results (as shown in Figure 11 in the appendix), which demonstrates that the propensity of a model to hallucinate is positively correlated between programming languages. The total hallucination rates for each evaluated model are presented in Figure 2. For a comprehensive breakdown of the results for each language, please refer to Appendix G. These results provide strong evidence that package hallucinations are a significant issue across all code generating LLMs.

### v2

@@ 3 Research Questions
Adversary Model and Assumptions. We assume an adversary that wants to execute a package confusion attack by leveraging package hallucinations in code generated by both closed-source and open-source code generation LLMs (see Figure 1). The target of the adversary are users who employ such LLMs for generating code. Here we assume that the LLM generates code that requires additional packages for the user to install, which the user does without sufficient verification of the recommended packages. In other words, the target users fully trust the LLMs to include only valid package names in the generated code. We assume that the adversary has access to the same set of LLMs for code generation as the target users, and is unable to modify or manipulate the model and model parameters of these LLMs (e.g., via re-training or fine-tuning) before they are used by the victims.
The adversary is able to determine a list of hallucinated packages generated by these LLMs (for example, by cross-referencing the package repository), and then is able to realize a package confusion attack by creating a package of the same name on the corresponding package repositories.
These newly created (and now publicly available) packages by the adversary could contain malicious code or functionality. Research has shown that installing open source Python or JavaScript packages allows arbitrary code execution to an attacker [57]. Previous work [35] has also established the viability of such an attack by publishing a hallucinated package to an open source repository and demonstrating that the package is actively downloaded and was incorporated into the dependency chains of other packages/code.
Research Questions. We now organize our investigation into five broad Research Questions (RQ).
RQ1: How prevalent are package hallucinations while generating Python and JavaScript code using LLMs? Our aim here is to thoroughly examine how often package hallucinations occur with both widely-used commercial and open-source LLMs when they generate Python and JavaScript code across various programming tasks.
RQ2: How are package hallucinations impacted by select model settings? Here our goal is to comprehensively analyze how training data and decoding strategies impact the package hallucinations produced by these code generation LLMs.
@@
RQ5: Is it possible to effectively mitigate package hallucinations using best practices in the literature and knowledge gained from earlier results? Through this RQ, we will investigate if code generating LLMs can be designed to reduce hallucinations with minimal compromise to code quality. In this direction, we will study if techniques such as retrieval augmented generation (RAG), self-detected feedback, decoding strategies, and supervised fine-tuning are effective hallucination reduction strategies.
@@ 4 Experiment Design
To address the RQs outlined above, we design several experiments to repeatedly prompt LLMs to generate code and then analyze the generated code. Our experimentation pipeline comprises of three distinct phases: (i) prompt dataset generation, (ii) code generation, and (iii) hallucination detection, each of which is described below:
@@ 4.1 Prompt Dataset
The experiments are designed to exhaustively test each LLM through a complete range of coding tasks. Existing datasets of benchmark coding prompts contain only a limited number of prompts (e.g. only 164 prompts for both EvalPlus [45] and HumanEval [8]) and lack diversity. Therefore, we create a new code prompt dataset for our experiment that contains both breadth and depth in terms of overall number of prompts and range of topics. Our goal was to develop a dataset that accurately and comprehensively represents coding tasks commonly requested by everyday users. To accomplish this, we employed two distinct approaches, as described below.
@@
Stack Overflow Dataset. To model the input prompts around real programmer questions, our first prompt dataset was created using Stack Overflow [70] questions across relevant programming topics and subject areas. Stack Overflow is a popular online question-and-answer service for software programmers and developers. To capture a wide range of topics, we utilized the “tag” feature of Stack Overflow, which allows users to label posts according to a subject matter. We included any tag that had over 5,000 questions and was also relevant to Python or JavaScript (the two programming languages we focus in this work, as detailed in Section 4.2). For each of the 240 manually selected tags that met this criteria (full list of tags is listed in Appendix H, we extracted the 20 most up-voted questions, resulting in 4,800 prompts (i.e., 4,800 prompts for Python and 4,800 prompts for JavaScript).
@@
As more recent data is less likely to be included in the pre-training data of LLMs, we are also interested in investigating the correlation between data recency (i.e., how recently the question was asked on Stack Overflow) and model hallucination rate. To enable such an analysis, we ran two queries on Stack Overflow; one that only captured the most popular questions in the selected tags from 2023 and another that captured the most popular questions for all years prior to 2023. By including the two different ranges of time, we effectively doubled the original number of prompts, for a total of 9,600 for each of the two languages.
Not all questions asked on Stack Overflow may involve coding or require code to answer the question. Rather than attempting to filter out such prompts during the code generation phase, which is non-trivial and error-prone, the LLM is asked to answer the question and only provide code if necessary. In the end, this may result in a slightly lesser number of usable LLM-generated code samples, but is more realistic as LLMs are expected to accommodate imperfect user inputs.
@@
LLM-generated Dataset. As a majority of the programming tasks require some library/package, our next idea was to use the package repositories themselves as a good representation of the full spectrum of coding topics. Our goal was to represent as many code generation tasks as possible in one comprehensive dataset. We take the 5,000 most popular Python and JavaScript packages (according to number of downloads) and scrape the official package description as listed on PyPI and npm, respectively. These descriptions are then input individually to the Llama-2 70B model with instructions to generate a coding prompt based on the package description (exact prompt available in Figure 13). This process worked remarkably well and generated roughly 4,800 prompts for Python and JavaScript each, resulting in two datasets of approximately the same size (some packages with no description or descriptions in a non-English language were discarded). Similar to the Stack Overflow dataset, we doubled the LLM-generated dataset for temporal analysis by dividing it into two segments: packages most downloaded in 2023 and packages most downloaded prior to 2023. When a package appears in both sets, we remove the package from the latter set to ensure no overlap. Removing duplicates from the latter dataset guarantees that the remaining packages will be those that have increased in popularity during the last year, capturing the desired signal. A truncated list from the LLM-generated dataset can be found in Appendix A.
@@ 4.2 Code Generation
Model Selection. For our experiments, we chose the top-ranked models from the EvalPlus leaderboard (as of January 2024) [45]. While creating our list, we disregarded the fine-tuned versions that were ranked below their corresponding foundational models, and only selected one fine-tuned version of the same foundational model of the same parameter size [45]. EvalPlus maintains a ranking of the top performing LLMs for code correctness according to a rigorous code synthesis evaluation framework. Our goal was to include a mix of top performing base models and a few of the best performing fine-tuned variants. We also included the GPT series of models (GPT-3.5, GPT-4, and GPT-4 Turbo) in our experiments, which currently hold the top rankings on the leaderboard. GPT models are widely considered as SOTA in terms of code generation models at the time of writing and add value to our experiments as representative commercial models. The models were not modified or altered in any way prior to testing; they are strictly “off-the-shelf.” Table 1 provides a complete list of the models that we tested in our experiments.
Table 1: Details of the models that were evaluated.
@@
Language Selection. In our experiments, we focus on two of the most popular programming languages, JavaScript and Python. These languages were chosen for their overall popularity (#1 and #2 according to GitHub’s 2023 Octoverse report [15]) and their reliance on open-source repositories for package management. Other popular programming languages like Java, C, or C++ do not rely on a centralized, open-source repository as Python and JavaScript do, which is a key component of this vulnerability. The open-source package repositories for these languages, npm and PyPI, represent ecosystems of 5.1 million and 573 thousand packages, respectively [72]. Of the 16 total models tested (see Table 1), fourteen were tested for both Python and JavaScript, while two fine-tuned Python-specific models, WizardCoder-Python and CodeLlama-Python, were only tested for Python.
Testing Environment. To create a fair and uniform testing environment, we utilize the text-generation-webui tool for LLMs, which supports most open-source LLMs and offers both a GUI and API for local and offline use [59]. It allows precise parameter control, ensuring consistent model evaluation during testing. We employed the GPTQ quantization method on these models, which reduces parameter precision to boost inference speed and lower memory use without significantly impacting performance. The GPTQ quantization method, specifically, uses a one-shot weight quantization method based on approximate second-order information that has a negligible effect on the accuracy of models, making it an ideal choice [13]. Additionally, quantized models better simulate the performance that a typical user can expect when running models on commercial grade hardware, making them more accessible and practical for everyday use.
For testing uniformity, we use the same parameters and quantization precision for all open-source models, which are summarized in Appendix C, along with the computing environment used. To generate code for our analysis, we query each of the LLMs (Table 1) with prompts from the two datasets using the Python requests packages to interact with the text-generation-webui API. The message to the API contains a system message with specific instructions for the model and the prompt itself. An overview of the process, including system messages used during each step, are detailed in Appendix B. This results in a total of 19,200 code samples per model (16 Python tests + 14 JavaScript tests * 19,200 = 576,000 total code samples), which are further analyzed to determine which packages are required to execute the generated code.
@@ 4.3 Detection Methodology and Heuristics
Detection of hallucinated packages from LLM outputs or code samples is non-trivial. In short, simply parsing the code for “import” or “require” is not useful as the arguments in those statements refer to modules and not packages. There is no way to definitively determine required packages from a code snippet alone. A detailed explanation of this problem can be found in Appendix G. To solve this problem, we employ the following three heuristics to determine/identify package names in the generated code:
@@
Heuristic 2. For the second heuristic, each generated code sample is used as an input into the same model that generated it. The model is then prompted for a list of packages that would be required to run the given code. Our intuition is to mimic an actual user/developer who is using LLMs for code generation. If the user gets an error due to an uninstalled package when attempting to execute the generated code, they could query the model for the correct package to install. We wanted to replicate this intuitive process to identify package names required by the generated code.
Heuristic 3. As the third heuristic, we re-use the original prompt used to generate the code sample as an input to the model and ask the model to output package names that would be required to accomplish this coding task. Similar to the previous heuristic, this process of extracting package names simulates another approach users would take to obtain package names from the model that generated the code, if the required packages were not mentioned in explicit “pip install” and “npm install” commands.
Once each model provides specific package names (through the three heuristics outlined above), we simply compare each package name to a master list of package names acquired from PyPI and npm, respectively (each list is as of 10 January, 2024). If a package name is not on the master list it is considered a hallucination. We acknowledge the possibility that the master list of packages obtained from the package repositories has already been contaminated with malicious hallucinated packages. It is not possible to guarantee that the master list actually represents the ground truth of valid packages, however, the presence of hallucinated packages already in the master list would actually produce fewer hallucinations, therefore our results represent a lower bound of hallucination rate.
@@ 5 Evaluation Results
In this section, we present the results of our experimental analysis related to RQ1 - RQ4. After using both Python and JavaScript for RQ1, for RQs 2 through 4, we focus our analysis only on the Python programming language, a subset of the original models tested, and randomly sampled subsets of our original datasets. Given the consistent results that we were able to obtain across both languages in RQ1, this narrowed scope of discussion does not compromise the generalizability of the conclusions and allows for a deeper analysis of package hallucinations in a controlled setting. We selected GPT-4 Turbo, GPT-3.5, CodeLlama 7B, and DeepSeek 6.7B for in-depth analysis, representing the best-performing and most popular open-source models.
@@ 5.1 Prevalence of Package Hallucinations (RQ1)
In our first experiment, our goal was to quantify the prevalence of package hallucinations across different models by generating and analyzing a large number of code samples. We conducted 30 tests (using 16 models for Python and 14 models for JavaScript, as outlined in Table 1) producing a combined 576,000 code samples using both the Stack Overflow and LLM-generated datasets (Section 4.1). Each code sample was evaluated for hallucinations according to the heuristics defined in Section 4.3, which include parsing the generated code and prompting the model for packages twice per code sample, for a total of 1,152,000 package prompts across all tests. To measure LLMs’ propensity to produce hallucinated packages during code generation we use the package hallucination rate metric, which can be expressed as a simple ratio of the number of hallucinated packages to the total number of recommended packages. The full set of hallucination results for all models, covering both Python and JavaScript, are presented in Appendix E.
@@
These 30 tests generated a total of 2.23 million packages in response to our prompts, of which 440,445 (19.7%) were determined to be hallucinations, including 205,474 unique non-existent packages (i.e. packages that do not exist in PyPI or npm repositories and were distinct entries in the hallucination count, irrespective of their multiple occurrences). Our results for GPT-3.5 (5.76%) and GPT-4 (4.05%) differ significantly from previous work on package hallucinations [35], which found hallucination rates 4-6 times greater (24.2% and 22.2%, respectively) for those specific models. GPT-series models were found 4 times less likely to generate hallucinated packages compared to open-source models, with a 5.2% hallucination rate compared to 21.7%. GPT-4 Turbo resulted in the lowest overall hallucination rate at 3.59%, while DeepSeek 1B had the best hallucination rate among the open-source models at 13.63%. Python code resulted in fewer hallucinations than JavaScript (15.8% on average compared to 21.3% for JavaScript). Despite the difference in the hallucination rate between the two languages, there is a linear relationship between the results (as shown in Figure 14 in the appendix), which demonstrates that the propensity of a model to hallucinate is positively correlated between programming languages. The total hallucination rates for each evaluated model are presented in Figure 2. For a comprehensive breakdown of the results for each language, please refer to Appendix E. These results provide strong evidence that package hallucinations are a pervasive issue across all code generating LLMs.
RQ1 Summary: Package hallucinations were found to be pervasive phenomenon across all 16 models tested. Commercial models hallucinated 4x less compared to open-source models. Python code resulted in a lower hallucination rate compared to JavaScript.